# Describe Where You Are: Improving Noise-Robustness for Speech Emotion Recognition with Text Description of the Environment

### v1

@@ Abstract
Speech emotion recognition (SER) systems often struggle in real-world environments, where ambient noise severely degrades their performance. This paper explores a novel approach that exploits prior knowledge of testing environments to maximize SER performance under noisy conditions. To address this task, we propose a text-guided, environment-aware training where an SER model is trained with contaminated speech samples and their paired noise description. We use a pre-trained text encoder to extract the text-based environment embedding and then fuse it to a transformer-based SER model during training and inference. We demonstrate the effectiveness of our approach through our experiment with the MSP-Podcast corpus and real-world additive noise samples collected from the Freesound repository. Our experiment indicates that the text-based environment descriptions processed by a large language model (LLM) produce representations that improve the noise-robustness of the SER system. In addition, our proposed approach with an LLM yields better performance than our environment-agnostic baselines, especially in low signal-to-noise ratio (SNR) conditions. When testing at -5dB SNR level, our proposed method shows better performance than our best baseline model by 31.8 % (arousal), 23.5% (dominance), and 9.5% (valence).
@@ 1 Introduction
Speech emotion recogntion (SER) systems have highly improved with the help of pre-trained speech representation models [1, 2, 3] and the creation of larger emotional speech databases [4, 5, 6, 7]. Recently, there has been increased interest in deploying SER systems in real-world applications, opening opportunities across many domains, such as digital assistants [8], health care applications [9], and security and defense. One important barrier in this direction is the degradation of SER performance in real-world environments caused by multiple types of non-stationary background noise [10].
@@
Several solutions have been proposed to improve the robustness of SER systems against acoustic noise. The solutions include data augmentation [11, 12], feature enhancement [13, 14], feature selection [15, 16], and domain adaptation approaches [17, 18]. Since transformer-based speech representation models have been successfully used in speech problems [1, 2, 3], many studies have also worked on increasing the noise robustness of SER systems built with pre-trained speech representation models [19, 20]. These approaches can increase the performance of transformer-based SER models in target noisy conditions. However, it is challenging to use these models in scenarios with multiple noisy environments since a transformer-based SER model requires important resources to adapt and store its parameters for each target environment. To address multiple noise types in a single SER model, Leem et al. [21] proposed environment-agnostic and -specific adapters. Their work showed that leveraging the prior knowledge of the testing condition is important for an SER model’s adaptation to multiple noisy environments.
@@
This paper focuses on how to effectively use the prior knowledge of a testing condition for an SER model that is adapted to multiple environments. The prior knowledge is used as a mechanism for zero-shot learning in new environments with types of noises not considered while training the models. It also provides the mechanism to indirectly identify similar environmental conditions during training (e.g., noise in a bus station and a train station). Exploring this problem, we investigate using text-based environment descriptions as the prior knowledge for a noise-robust SER system. Using natural language prompts during training has shown potential in image classification [22], sound event classification [23], and several speech processing downstream tasks, including keyword spotting, and speaker counting [24]. Natural language supervision is also applicable to SER tasks [25, 26]. All these studies indicate that exploiting text information is a promising strategy to SER systems. We propose a text-guided environment-aware training (TG-EAT) strategy to improve the noise robustness of an SER model with text descriptions. We focus on the prediction of arousal (calm to active), valence (negative to positive), and dominance (weak to strong). TG-EAT uses noisy speech and its text-based environmental description to adapt the SER model. We use a pre-trained text encoder to extract the representation of text-based environment descriptions. This representation is combined with a transformer-based SER model. During adaptation, the SER model learns appropriate denoising functions with respect to the given environment description. During inference, we only need to change the template sentence to guide the SER model with testing environment information. We expect that the pre-trained text encoder can capture similar semantic information from environmental conditions included in the train set, allowing zero-shot environment learning for the SER model. This approach is expected to generalize the SER performance when tested in environmental conditions that are not included in the training process.
@@
Our experiment with the MSP-Podcast corpus shows that using text description of the testing environment can highly improve the SER performance, especially with large language model (LLM). In the -5dB signal-to-noise ratio (SNR) condition, our method improves the original SER model built with a self-supervised learning (SSL) representation by 163.6% for arousal, 200.0% for dominance, and 91.6% for valence. When we compare the proposed SER model with our best baseline, we observe improvements of 31.8 % for arousal, 23.5% for dominance, and 9.5% for valence (-5dB SNR level). With the text encoder from CLAP, pre-trained with paired audio, the SER model can achieve the best performance in the low SNR condition. Compared with freezing the text encoder, the fine-tuning approach improves performance by 72.2% for arousal, 91.6% for dominance, and 21.0% for valence under the -5dB SNR condition. Our solution is highly applicable to SER systems deployed in real-world applications. For example, systems can infer the testing environment from the global positioning system (GPS) information by using geological information service (GIS) mashups, such as OpenStreetMap [27]. The main contributions of this study are:
  • We explore using text embedding for an SER model to increase noise robustness in unseen conditions by explicitly leveraging the environment information.
@@ 2.1 Speech Emotion Recognition under Noisy Environments
Increasing the noise robustness of an SER system is an essential task when deploying it in real-world applications. Previous studies have mainly focused on improving acoustic features for the SER model. Triantafyllopoulos et al. [14] proposed to enhance noisy low-level descriptors (LLDs) for an SER model by using a convolutional neural network (CNN) with residual blocks. Pandharipande et al. [28]proposed to discard noisy frames to increase the noise robustness of an SER model by using a voice activity detection module. Leem et al. [29] proposed to select noise-robust LLDs by addressing the performance and robustness of each single LLD.
@@
More recently, SER studies have mainly focused on using transformer-based speech representation models [30, 31, 32, 33, 34, 35], including Wav2Vec2.0 [1], HuBERT [2], and WavLM [3]. Such models have shown higher robustness against the small perturbation on the input speech than the traditional SER model with a Mel-spectrogram [32]. Despite this trend, they still show performance differences from the ones tested in a clean environment. For this reason, studies are currently exploring strategies to improve the noise robustness of the pre-trained speech representation model. A common approach to address this issue is noise-aware training, where the clean training set is augmented with the noise sound during environment adaptation. Mitra et al. [19] demonstrated that training a HuBERT-based SER model with noisy speech can highly improve the performance in low signal-to-noise ratio (SNR) conditions. Leem et al. [20] proposed a contrastive teacher-student learning strategy to address the catastrophic forgetting issue when training a fine-tuned SER model with noisy speech. Wu et al. [12] proposed to dynamically change the distortion level of the augmented speech during adaptation based on the distortion metrics.
@@
The aforementioned methods focused on increasing the SER model’s robustness against a single target environment. They might not be the optimal solution for an SER model deployed on a real-world application since it is highly likely that this system will encounter multiple types of environmental noises. We focus on adapting a single transformer-based SER model to multiple noisy environments to efficiently deal with multiple types of environments. To address this issue, Leem et al. [21] proposed to adapt the transformer-based SER model to multiple types of noises with skip connection adapters. They not only trained the SER model with multiple environments but also focused on leveraging the environmental information of the testing conditions to improve SER performance under noisy conditions. The results showed that using the environment-agnostic and -specific adapters with respect to the testing condition can improve the SER performance under noisy conditions. Such prior knowledge could be achieved using domain knowledge or global positioning system (GPS) information. Their result showed that using environmental information during inference is important for a SER model to perform well under noisy conditions. This work indicates that leveraging the prior knowledge of the testing condition is also important for a noise-robust SER model, as well as training it with multiple types of noises. This is beneficial for an SER model deployed on real-world applications where the system can exploit the domain knowledge of the testing environment and the global positioning system (GPS) information.
@@ 2.2 Text-Guided Training
As we discussed in Section 2.1, exploiting environmental information can improve SER performance in a noisy environment. This paper mainly focuses on using text prompts to infuse environmental information into an SER model. Using natural language prompts does not require the recognition model to use a fixed set of predetermined labels during training. Contrastive language-image pre-training (CLIP) is a good example of this approach [22]. It consists of an image encoder and a text encoder, trained with pairs of images and their corresponding text descriptions. These encoders are trained in a contrastive learning manner, which maximizes the similarity of both representations if the image and the description are paired and minimizes the similarity if they are unpaired. After training, these encoders can perform zero-shot classification by checking the similarity between the given image and the candidate prompts. The study of Radford et al. [22] used the following prompt template: “A photo of a {label}. They calculate the similarity between the representation from the given image and the representations from the prompts with different {label}, selecting the {label} that shows the maximum similarity.
The contrastive pre-training strategy with natural language supervision is also successful in universal audio and speech processing. Wu et al. [23] demonstrated that pre-training audio and text encoder with natural language guidance could improve audio classification performance. The study of Elizalde et al. [24] showed that such natural language guidance can improve speech processing tasks, including keyword spotting, speaker counting, and SER tasks.
Previous studies have found that natural language supervision can apply to SER tasks. Stanley et al. [25] used word embeddings to encode emotional labels for SER model. Gong et al. [26] used large language model (LLM) to infer weak emotion labels for unlabeled data for weakly-supervised learning of an SER model. All these findings have shown that exploiting text information is highly applicable to SER systems. To the best knowledge of the authors, the use of natural language supervision to address SER robustness against unknown noisy environments is a novel research direction.
@@ 3 Proposed Method
This paper proposes text-guided environment-aware training (TG-EAT), which leverages environmental information to improve an SER model in noisy conditions. Figure 1 illustrates our proposed TG-EAT framework, which uses a pair of noisy speech and its corresponding environmental description. The text embedding extracted from the environmental description is combined with the acoustic representation in the SER model, allowing it to denoise the representation for the given environmental description.
@@
The key contribution of this study is how we use the text description from the target environment. We used prompts to generate the text description where the target environment is changed. As a preliminary experiment, we tested different prompts to describe the target environment such as “The type of background noise is {environment},” or “The input is recorded with a sound of {environment}.” We change {environment} in the prompts according to the target environment during training and testing. We found that all the prompts showed similar emotion recognition performance for all the attributes. Therefore, we consistently use the following prompt in this study: “This speech is recorded in {environment}.” We extract the text-based environment embedding from this text description using a pre-trained text encoder. We test two different text representations: contrastive learning (CL)-based representation and LLM-based representation. For the CL-based representation, we use the text encoder pre-trained with the contrastive language audio pre-training (CLAP) strategy [23, 24]. CLAP consists of an audio encoder and a text encoder. It uses a pair of acoustic events and their text description during pre-training (e.g., bird chirping sound with the description, “Bird is chirping in the given audio”). With these audio-text pairs, the training objective is to maximize the similarity between the audio and text representation if they are from the same pair and minimize it if they are from a different pair. Since CLAP uses an audio-text pair during pre-training, we assume that its text encoder can generate an appropriate representation from the given environment description coherent with the target acoustic condition. This paper uses the pre-trained text encoder from the unfused CLAP model proposed in the study of Wu et al. [23]. For the LLM-based representation, we use the encoder from the pre-trained RoBERTa model [36]. RoBERTa is pre-trained with masked language modeling (MLM) and next sentence prediction (NSP) tasks. RoBERTa has shown good performance in various benchmarks for evaluating natural language understanding systems, such as GLUE [37]. Although it is not pre-trained with audio data, we assume that its encoder can extract enriched semantic information from the given prompt. We use RoBERTa-large, which has 24 transformer layers. For each text encoder, we use the same tokenizer used in its pre-training to tokenize the text description of the environment. We extract token-level text embeddings from the tokenized prompt and then apply average pooling, resulting in a single representation vector for each prompt.
@@
After the environmental representation is obtained, the next step is to introduce this information into the model. We mainly focus on a transformer-based SER model, which has shown good performance in SER tasks [38, 32]. An important task is to fine-tune the model with clean and emotional speech data. We first fine-tune the SER model with clean speech to maximize the concordance correlation coefficient (CCC) between the predicted and the ground-truth emotional attribute scores of arousal, dominance, and valence. After fine-tuning with clean speech, the SER model is continuously updated with the training set contaminated with multiple types of noise and their corresponding text description. We insert the text representation from the given environment description into the fine-tuned transformer-based SER model. We achieve this goal by combining the text embedding with the acoustic representation, which is the output of the convolutional encoder. We apply trainable linear projection to the text embedding to match its dimension to the acoustic representation embeddings. We concatenate the projected text embedding to the acoustic representation embeddings along the time axis, then feed them into the transformer encoder. We update the transformer encoder and the downstream head with the concatenated embeddings. We use the same training objective as the one used when training with clean speech. From this framework, we want to evaluate if the SER model can learn the denoising function given a noisy acoustic representation with its text embedding.

### v2

@@ Abstract
Speech emotion recognition (SER) systems often struggle in real-world environments, where ambient noise severely degrades their performance. This paper explores a novel approach that exploits prior knowledge of testing environments to maximize SER performance under noisy conditions. To address this task, we propose a text-guided, environment-aware training where an SER model is trained with contaminated speech samples and their paired noise description. We use a pre-trained text encoder to extract the text-based environment embedding and then fuse it to a transformer-based SER model during training and inference. We demonstrate the effectiveness of our approach through our experiment with the MSP-Podcast corpus and real-world additive noise samples collected from the Freesound and DEMAND repositories. Our experiment indicates that the text-based environment descriptions processed by a large language model (LLM) produce representations that improve the noise-robustness of the SER system. With a contrastive learning (CL)-based representation, our proposed method can be improved by jointly fine-tuning the text encoder with the emotion recognition model. Under the -5dB signal-to-noise ratio (SNR) level, fine-tuning the text encoder improves our CL-based representation method by 76.4% (arousal), 100.0% (dominance), and 27.7% (valence).
@@ 1 Introduction
Speech emotion recogntion (SER) systems have highly improved with the help of pre-trained speech representation models [Baevski_2020, Hsu_2021, Chen_2022] and the creation of larger emotional speech databases [Martinez-Lucas_2020, Kondratenko_2023, Lotfian_2019_3, Upadhyay_2023_2]. Recently, there has been increased interest in deploying SER systems in real-world applications, opening opportunities across many domains, such as digital assistants [Chatterjee_2021], health care applications [Fulford_2021], and security and defense. One important barrier in this direction is the degradation of SER performance in real-world environments caused by multiple types of non-stationary background noise [Lee_2021].
@@
Several solutions have been proposed to improve the robustness of SER systems against acoustic noise. The solutions include data augmentation [lakomkin_2018, Wu_2023_2, Ranjan_2024], feature enhancement [juszkiewicz_2014, Triantafyllopoulos_2019], feature selection [Schuller_2006, Leem_2022], and domain adaptation approaches [Wilf_2021, Leem_2021]. Since transformer-based speech representation models have been successfully used in speech problems [Baevski_2020, Hsu_2021, Chen_2022], many studies have also worked on increasing the noise robustness of SER systems built with pre-trained speech representation models [Mitra_2023, Leem_2023]. These approaches can increase the performance of transformer-based SER models in target noisy conditions. However, it is challenging to use these models in scenarios with multiple noisy environments since a transformer-based SER model requires important resources to adapt and store its parameters for each target environment. To address multiple noise types in a single SER model, Leem et al. [Leem_2023_2] proposed environment-agnostic and -specific adapters. Their work showed that leveraging the prior knowledge of the testing condition is important for an SER model’s adaptation to multiple noisy environments.
@@
This paper explores which form of prior knowledge allows an SER model to effectively adapt to multiple unseen environments. Rather than aiming to cover all the environments, our system trained the model to be conditioned by a text embedding describing the environment, which project the unseen condition into the ones that are the closest to the seen environments. With this strategy, the prior knowledge is used as a mechanism for zero-shot learning in new environments with types of noises not considered while training the models. It also provides the mechanism to indirectly identify similar environmental conditions during training (e.g., noise in a bus station and a train station). Exploring this problem, we investigate using text-based environment descriptions as the prior knowledge for a noise-robust SER system. Using natural language prompts during training has shown potential in image classification [Radford_2021], sound event classification [Wu_2023], and several speech processing downstream tasks, including keyword spotting, and speaker counting [Elizalde_2023]. Natural language supervision is also applicable to SER tasks [Stanley_2023, Gong_2023]. All these studies indicate that exploiting text information is a promising strategy for SER systems. We propose a text-guided environment-aware training (TG-EAT) strategy to improve the noise robustness of an SER model with text descriptions. We focus on the prediction of arousal (calm to active), valence (negative to positive), and dominance (weak to strong). TG-EAT uses noisy speech and its text-based environmental description to adapt the SER model. We use a pre-trained text encoder to extract the representation of text-based environment descriptions. This representation is combined with a transformer-based SER model. During adaptation, the SER model learns appropriate denoising functions with respect to the given environment description. During inference, we only need to change the template sentence to guide the SER model with testing environment information. We expect that the pre-trained text encoder can capture similar semantic information from environmental conditions included in the train set, allowing zero-shot environment learning for the SER model. This approach is expected to generalize the SER performance when tested in environmental conditions that are not included in the training process.
@@
Our experiment with the MSP-Podcast corpus shows that using text descriptions of the testing environment can highly improve the SER performance, especially with large language model (LLM). In the -5 dB signal-to-noise ratio (SNR) condition, our method improves the original SER model built with a self-supervised learning (SSL) representation by 7.6% for arousal, 8.3% for dominance, and 45.4% for valence. When we compare the proposed SER model with the DAT baseline, we observe improvements of 16.6% for arousal, 18.1% for dominance, and 23.0% for valence (-5 dB SNR level). With the text encoder from CLAP, pre-trained with paired audio, the SER model can achieve the best performance in the low SNR condition. Compared to freezing the text encoder, the fine-tuning approach improves performance by 76.4% for arousal, 100.0% for dominance, and 27.7% for valence under the -5 dB SNR condition. Our solution is highly applicable to SER systems deployed in real-world applications. For example, systems can infer the testing environment from a global positioning system (GPS) by using geological information service (GIS) mashups, such as OpenStreetMap [Vargas-Munoz_2021]. The main contributions of this study are:
  • We explore using text embedding for an SER model to increase noise robustness in unseen conditions by explicitly leveraging the environment information. Our method provides a unique advantage by enabling a single model to adapt to multiple noise conditions using text embeddings rather than requiring multiple context-specific expert models. This is particularly beneficial for transformer-based architectures, which demand significant computational resources while delivering SOTA performance for SER. By leveraging text-described target environment information, we maximize performance without the overhead of maintaining multiple models.
@@ 2.1 Speech Emotion Recognition under Noisy Environments
Increasing the noise robustness of an SER system is an essential task when deploying it in real-world applications. Previous studies have mainly focused on improving acoustic features for the SER model. Triantafyllopoulos et al. [Triantafyllopoulos_2019] proposed to enhance noisy waveforms before extracting the SER features. The enhancement models used convolutional neural network (CNN) with residual blocks. Pandharipande et al. [Pandharipande_2018] proposed to discard noisy frames to increase the noise robustness of an SER model by using a voice activity detection module. Leem et al. [Leem_2024] proposed to select noise-robust LLDs by addressing the performance and robustness of each single LLD.
@@
More recently, SER studies have mainly focused on using transformer-based speech representation models [Goncalves_2024, Pepino_2021, Wagner_2023, Mote_2024, Goncalves_2024_2, Upadhyay_2024], including Wav2Vec2.0 [Baevski_2020], HuBERT [Hsu_2021], and WavLM [Chen_2022]. Such models have shown higher robustness against small perturbations on the input speech than the traditional SER model with a Mel-spectrogram [Wagner_2023]. Despite this trend, they still show performance differences from the ones tested in a clean environment. For this reason, studies are currently exploring strategies to improve the noise robustness of the pre-trained speech representation model. A common approach to address this issue is noise-aware training, where the clean training set is augmented with the noise sound during environment adaptation. Mitra et al. [Mitra_2023] demonstrated that training a HuBERT-based SER model with noisy speech can highly improve the performance in low SNR conditions. Leem et al. [Leem_2023] proposed a contrastive teacher-student learning strategy to address the catastrophic forgetting issue when training a fine-tuned SER model with noisy speech. Wu et al. [Wu_2023_2] proposed to dynamically change the distortion level of the augmented speech during adaptation based on the distortion metrics.
@@
The aforementioned methods focused on increasing the SER model’s robustness against a single target environment. They might not be the optimal solution for an SER model deployed on a real-world application since it is highly likely that this system will encounter multiple types of environmental noises. We focus on adapting a single transformer-based SER model to multiple noisy environments to efficiently deal with multiple types of environments. To address this issue, Leem et al. [Leem_2023_2] proposed to adapt the transformer-based SER model to multiple types of noises with skip connection adapters. They not only trained the SER model with multiple environments but also focused on leveraging the environmental information of the testing conditions to improve SER performance under noisy conditions. The results showed that using the environment-agnostic and -specific adapters with respect to the testing condition can improve the SER performance under noisy conditions. Such prior knowledge could be achieved using domain knowledge or GPS information. Their result showed that using environmental information during inference is important for a SER model to perform well under noisy conditions. This work indicates that leveraging the prior knowledge of the testing condition is also important for a noise-robust SER model, as well as training it with multiple types of noises. This is beneficial for an SER model deployed on real-world applications where the system can exploit the domain knowledge of the testing environment and the GPS information.
@@ 2.2 Text-Guided Training
As we discussed in Section 2.1, exploiting environmental information can improve SER performance in a noisy environment. This paper mainly focuses on using text prompts to infuse environmental information into an SER model. Using natural language prompts does not require the recognition model to use a fixed set of predetermined labels during training. Contrastive language-image pre-training (CLIP) is a good example of this approach [Radford_2021]. It consists of an image encoder and a text encoder, trained with pairs of images and their corresponding text descriptions. These encoders are trained in a contrastive learning manner, which maximizes the similarity of both representations if the image and the description are paired and minimizes the similarity if they are unpaired. After training, these encoders can perform zero-shot classification by checking the similarity between the given image and the candidate prompts. The study of Radford et al. [Radford_2021] used the following prompt template: “A photo of a {label}. They calculate the similarity between the representation from the given image and the representations from the prompts with different {label}, selecting the {label} that shows the maximum similarity.
The contrastive pre-training strategy with natural language supervision is also successful in universal audio and speech processing. Wu et al. [Wu_2023] demonstrated that pre-training audio and text encoder with natural language guidance could improve audio classification performance. The study of Elizalde et al. [Elizalde_2023] showed that such natural language guidance can improve speech processing tasks, including keyword spotting, speaker counting, and SER tasks.
Previous studies have found that natural language supervision can apply to SER tasks. Stanley et al. [Stanley_2023] used word embeddings to encode emotional labels for SER model. Gong et al. [Gong_2023] used LLM to infer weak emotion labels for unlabeled data for weakly-supervised learning of an SER model. All these findings have shown that exploiting text information is highly applicable to SER systems. To the best knowledge of the authors, the use of natural language supervision to address SER robustness against unknown noisy environments is a novel research direction.
@@ 3 Proposed Method
This paper proposes text-guided environment-aware training (TG-EAT), which leverages environmental information to improve an SER model in noisy conditions. Figure 1 illustrates our proposed TG-EAT framework, which uses a pair of noisy speech and its corresponding environmental description. The text embedding extracted from the environmental description is combined with the acoustic representation in the SER model, allowing it to improve the representation for the given environmental description.
@@
The key contribution of this study is how we use the text description from the target environment. We used prompts to generate the text description where the target environment is changed. As a preliminary experiment, we tested different prompts to describe the target environment such as “The type of background noise is {environment},” or “The input is recorded with a sound of {environment}.” We change {environment} in the prompts according to the target environment during training and testing. We found that all the prompts showed similar emotion recognition performance for all the attributes. Therefore, we consistently use the following prompt in this study: “This speech is recorded in {environment}.” We extract the text-based environment embedding from this text description using a pre-trained text encoder. We test two different text representations: contrastive learning (CL)-based representation and LLM-based representation. For the CL-based representation, we use the text encoder pre-trained with the contrastive language audio pre-training (CLAP) strategy [Wu_2023, Elizalde_2023]. CLAP consists of an audio encoder and a text encoder. It uses a pair of acoustic events and their text description during pre-training (e.g., bird chirping sound with the description, “Bird is chirping in the given audio”). With these audio-text pairs, the training objective is to maximize the similarity between the audio and text representation if they are from the same pair and minimize it if they are from a different pair. Since CLAP uses an audio-text pair during pre-training, we assume that its text encoder can generate an appropriate representation from the given environment description coherent with the target acoustic condition. This paper uses the pre-trained text encoder from the unfused CLAP model proposed in the study of Wu et al. [Wu_2023]. We take a 768-dimensional latent embedding from the text encoder, using it as our text-based environmental embedding. For the LLM-based representation, we use the encoder from the pre-trained RoBERTa model [Liu_2019_2]. RoBERTa is pre-trained with masked language modeling (MLM) and next sentence prediction (NSP) tasks. RoBERTa has shown good performance in various benchmarks for evaluating natural language understanding systems, such as GLUE [Wang_2019_4]. Although it is not pre-trained with audio data, we assume that its encoder can extract enriched semantic information from the given prompt. We use RoBERTa-large, which has 24 transformer layers. For each text encoder, we use the same tokenizer used in its pre-training to tokenize the text description of the environment. We extract token-level text embeddings from the tokenized prompt and then apply average pooling, resulting in a 1024-dimensional single representation vector for each prompt.
@@
After the environmental representation is obtained, the next step is to introduce this information into the model. We mainly focus on a transformer-based SER model, which has shown good performance in SER tasks [Keesing_2021, Wagner_2023]. An important task is to fine-tune the model with clean and emotional speech data. We first fine-tune the SER model with clean speech to maximize the concordance correlation coefficient (CCC) between the predicted and the ground-truth emotional attribute scores of arousal, dominance, and valence. After fine-tuning with clean speech, the SER model is continuously updated with the training set contaminated with multiple types of noise and their corresponding text description. We insert the text representation from the given environment description into the fine-tuned transformer-based SER model. We achieve this goal by combining the text embedding with the acoustic representation, which is the output of the convolutional encoder. We apply trainable linear projection to the text embedding to match its dimension to the acoustic representation embeddings. We concatenate the projected text embedding to the acoustic representation embeddings along the time axis, then feed them into the transformer encoder. We choose this approach to allow the self-attention module in the transformer encoder to attend to the text embedding to all acoustic representation embeddings. Previous studies have proposed alternative approaches to add text embeddings into a prediction model [Perez_2018, Alayrac_2022], but we leave this research direction as our future investigation to further improve its performance. We update the transformer encoder and the downstream head with the concatenated embeddings. We use the same training objective as the one used when training with clean speech. From this framework, we want to evaluate if the SER model can learn the denoising function given a noisy acoustic representation with its text embedding.